OpenStudy (anonymous):

Please help write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirds x + 8 y = four thirds x + 3 y = -three fourths x + 8 y = -three fourths x + 3

OpenStudy (a_clan):

Two parallel lines have same slope. Hence, we find the slope of given line. Equation of line in slope-intercept form : y = mx+c We get: y= (4/3)x + 4 . Hence, slope = 4/3 which is also the slope of any line which is parallel to the given line. Our required line is : y= (4/3)x + C Now put the given point (-3, 4), through which this line passes, in the equation. 4 = (4/3) * -3 + C This gives C=8 Hence required equation is y = (4/3)x + 8